OpenStudy (anonymous):

What is the correct simplification of 2a^0?

OpenStudy (anonymous):

Anything raised to the power 0 is 1.

OpenStudy (anonymous):

So is it two or one?

OpenStudy (anonymous):

that depends on whether you meant (2a)^0 or 2 (a)^0

OpenStudy (anonymous):

There are no parentheses, it's just 2a^0.

OpenStudy (anonymous):

So the 2 isnt'raised to the power 0? In that case it's 2: 2a^0=2*1=2

OpenStudy (anonymous):

Okay, thanks!